# BabyVLM: Data-Efficient Pretraining of VLMs Inspired by Infant Learningthanks: Project website: shawnking98.github.io/BabyVLM

### v1

@@ 4.3 Assessing Language Bias in VTWT
4.4 Fine-Grained Analysis of Compositional Reasoning
@@
Type of Difference CVCL-ori CVCL-aug CVCL-aug-random
@@
Upper Bound Models LLaVA-v1.5-7B 0.7299 0.8300 0.6327 0.6273 0.8570
LLaVA-v1.5-7B-ft 0.7205 0.8032 0.5992 0.4941 0.6300
CLIP-large N/A N/A 0.5638 0.2397 0.7172
Baby Models BabyLLaVA 0.6772 0.5903 0.5214 0.2312 0.3907
CVCL N/A N/A 0.5221 0.1600 0.3993
Random Guess - 0.5000 0.5000 0.5000 0.1250 0.3750
@@
Additionally, models trained exclusively on developmentally plausible data (CVCL-ori and CVCL-aug) exhibit distinct performance patterns. For single-component differences (the first three rows of Table 6), both models achieve their highest performance on noun differences and perform worse on verb and adjective differences (e.g., 76% vs. 65% and 60% respectively for CVCL-aug). This result aligns with linguistic development findings from [6, 37], which suggest that early-stage language learners use nouns at least twice as often as verbs and are also slower to acquire adjectives. However, similar phenomenon is not observed from CVCL-aug-random.
This alignment reinforces that our targeted synthetic data transformations effectively facilitate more robust baby-like representations — a central objective identified in our introduction.
4.5 Evaluating Developmental Appropriateness
A primary aim of our approach is ensuring baby models align with the cognitive and linguistic limitations of early-stage learners. To empirically validate this property, we explicitly assess baby models on tasks that exceed typical infant-level developmental capacities, such as advanced visual reasoning (Winoground) and general-purpose tasks (VQA and BLiMP). As shown in Table 7, baby models (e.g., BabyLLaVA, CVCL) perform significantly below upper-bound models, affirming their constrained generalization capabilities. This limitation ensures developmental authenticity, preventing baby models from inadvertently solving complex tasks beyond its intended cognitive stage.
4.6 Out-of-Domain Generalization
While our primary goal is effective training within the SAYCam domain, assessing how models generalize to standard out-of-domain benchmarks remains insightful. Notably, as reinforced in the previous section, models intentionally designed within developmental constraints, such as BabyLLaVA and CVCL, show limited generalization to standard out-of-domain benchmarks like DevBench and VQA (Table 8). These results suggest that appropriate developmental modeling naturally constrains generalization, a desirable property confirming realistic cognitive boundaries rather than a limitation to be overcome.
@@ 5 Conclusion and Future Work
We construct Baby Winoground using the 967 test samples from the Visual Two-Word Test (VTWT). Our goal is to modify the original image from VTWT such that the modified image is associated exclusively with the negative phrase while preserving most of the original content. To achieve this, we leverage the search and replace functionality of Stability AI’s Stable Image Ultra model as our image-editing tool. This process requires two prompts:
@@
We train and evaluate two variants of the CVCL model from [50] (CVCL-aug & CVCL-aug-random). CVCL-aug is trained on our filtered SAYCam dataset and our transferred dataset (section 3), both of which contain approximately 67k image-caption pairs. Similarly, CVCL-aug-random is trained on our filtered SAYCam dataset plus a randomly sampled unprocessed subset of approximately 67k image-caption pairs from LLaVA’s pretraining dataset [22]. We train both variants on a single A100 GPU for 12 hours each using the default hyperparameters specified in the supplemental info of [50]. For evaluation, we use the model checkpoint from the last training epoch for each variant.
Out-of-Domain Ablation Study.
We also evaluate both our CVCL and BabyLLaVA model variants on several out-of-domain benchmarks, including general purpose benchmarks like VQA, and developmental benchmarks such as DevBench. We make several observations. First, we see that all model variants perform around random chance on Winoground, indicating that none of the models achieve robust compositional reasoning ability. For VQA and DevBench, however, both CVCL and BabyLLaVA variants trained on our transferred dataset (CVCL-aug & BabyLLaVA-aug) achieve superior performance, reinforcing the value of our developmentally adapted general-domain data. In addition, even the weakest BabyLLaVA model outperform all the CVCL variants on VQA, indicating the advanced reasoning ability of generative VLMs over discriminative VLMs. Finally, the modest performance across all model variants and tasks reinforces one of our main results that appropriate developmental modeling naturally constrains generalization.
@@
CVCL-ori N/A N/A 0.5221 0.1600 0.3993
CVCL-aug N/A N/A 0.4714 0.1641 0.6086
CVCL-aug-random N/A N/A 0.4935 0.1173 0.5198
BabyLLaVA-ori 0.6772 0.5903 0.5214 0.2312 0.3907
BabyLLaVA-aug 0.6646 0.5061 0.5455 0.4064 0.5303
BabyLLaVA-aug-random 0.6746 0.4778 0.5335 0.3659 0.4722

### v2

@@ 4.3 Assessing Language Bias in VTWT
4.4 Investigating Compositional Reasoning
@@
Type of Difference CVCL-filtered CVCL-filtered-aug CVCL-filtered-random
@@
Additionally, models trained exclusively on developmentally plausible data (CVCL-filtered and CVCL-filtered-aug) exhibit distinct performance patterns. For single-component differences (the first three rows of Table 6), both models achieve their highest performance on noun differences and perform worse on verb and adjective differences (e.g., 76% vs. 65% and 60% respectively for CVCL-filtered-aug). This result aligns with linguistic development findings from [6, 43], which suggest that early-stage language learners use nouns at least twice as often as verbs and are also slower to acquire adjectives. However, similar phenomenon is not observed from CVCL-filtered-random.
The alignment between our empirical results and developmental psychology literature reinforces that our targeted synthetic data transformations effectively facilitate more robust baby-like representations — a central objective identified in our introduction.
@@ 5 Conclusion and Future Work
We construct Baby Winoground using the 967 test samples from the Visual Two-Word Test (VTWT). Our goal is to modify the original image from VTWT such that the modified image is associated exclusively with the negative phrase while preserving most of the original content. To achieve this, we leverage the search and replace functionality of Stability AI’s Stable Image Ultra model [12] as our image-editing tool. This process requires two prompts:
@@
We train and evaluate two variants of the CVCL model from [57] (CVCL-filtered-aug & CVCL-filtered-random). CVCL-filtered-aug is trained on our filtered SAYCam dataset and our transferred dataset (section 3), both of which contain approximately 67k image-caption pairs. Similarly, CVCL-filtered-random is trained on our filtered SAYCam dataset plus a randomly sampled unprocessed subset of approximately 67k image-caption pairs from LLaVA’s pretraining dataset [28]. We train both variants on a single A100 GPU for 12 hours each using the default hyperparameters specified in the supplemental info of [57]. For evaluation, we use the model checkpoint from the last training epoch for each variant.
Out-of-Domain Generalization.
A primary aim of our approach is to ensure baby models align with the cognitive and linguistic limitations of early-stage learners. To empirically validate this property, we explicitly assess baby models on tasks that exceed typical infant-level developmental capacities, such as advanced visual reasoning (Winoground) and general-purpose tasks (VQA and BLiMP). As shown in Table 7, baby models (e.g., BabyLLaVA, CVCL) perform significantly below upper-bound models, affirming their constrained generalization capabilities. This limitation ensures developmental authenticity, preventing baby models from inadvertently solving complex tasks beyond their intended cognitive stage.
Upper Bound Models LLaVA-v1.5-7B 0.7299 0.8300 0.6327 0.6273 0.8570
LLaVA-v1.5-7B-ft 0.7205 0.8032 0.5992 0.4941 0.6300
CLIP-large N/A N/A 0.5638 0.2397 0.7172
Baby Models BabyLLaVA-GPT2 0.6772 0.5903 0.5214 0.2312 0.3907
BabyLLaVA-Llama 0.5095 0.5218 0.5094 0.1477 0.3082
CVCL N/A N/A 0.5221 0.1600 0.3993
Random Guess - 0.5000 0.5000 0.5000 0.1250 0.3750
Out-of-Domain Tasks Ablation Study.
We also evaluate both our CVCL and BabyLLaVA model variants on several out-of-domain benchmarks, including general purpose benchmarks like VQA, and developmental benchmarks such as DevBench. We make several observations. First, we see that all model variants perform around random chance on Winoground, indicating that none of the models achieve robust compositional reasoning ability. For VQA and DevBench, however, both CVCL and BabyLLaVA variants trained on our transferred dataset (CVCL-filtered-aug & BabyLLaVA-filtered-aug) achieve superior performance, reinforcing the value of our developmentally adapted general-domain data. In addition, even the weakest BabyLLaVA model outperform all the CVCL variants on VQA, indicating the advanced reasoning ability of generative VLMs over discriminative VLMs. Finally, the modest performance across all model variants and tasks reinforces one of our main results that appropriate developmental modeling naturally constrains generalization.
@@
CVCL-filtered N/A N/A 0.5221 0.1600 0.3993
CVCL-filtered-aug N/A N/A 0.4714 0.1641 0.6086
CVCL-filtered-random N/A N/A 0.4935 0.1173 0.5198
BabyLLaVA-filtered 0.6772 0.5903 0.5214 0.2312 0.3907
BabyLLaVA-filtered-aug 0.6646 0.5061 0.5455 0.4064 0.5303
BabyLLaVA-filtered-random 0.6746 0.4778 0.5335 0.3659 0.4722
@@
Transferred Data Efficiency Ablation Study.
To further investigate the extra data efficiency brought by our introduced transferred dataset, we subsample filtered-aug and filtered-random datasets by different fractions, then perform model training with identical training steps, and test on our in-domain benchmarks. Figure 12 shows the result. The filtered-aug curve rises more steeply, and using only 25–50% of filtered-aug already equals the full filtered-random on in-domain tasks, confirming substantial sample-efficiency gains.
Figure 12: Performance on different fractions of datasets.